suppose that the current spot exchange rate is €0.815/$ and the three month forward exchange rate is €0.815/$. the three month interest rate is 6.00 percent per annum in the United States and 5.40 percent per annum in France . assume that you can borrow up to $1,000,000 or €30,000. show how to realize a certain profit via covered interest arbitrage, assuming that you want to realize profit in terms of U.S dollars. also determine the size of your arbitrage profit.
Answer #1
The premium/(discount) on the euro = 0, as the spot and forward
rates are the same.
The interest rate differnential = 6.00-54.40 = 0.60%.
CIA is possible and as the interest rate differential is more it would
be beneficial to invest in the currency having higher interest rate.
Strategy would be to borrow in Euro and to invest in $.
Borrow Euro 30000 to finally repay = 30000*(1+0.054/4) = €     30,405.00
Convert the Euros to $ to get 30000/0.815 = $     36,809.82
Invest the dollars for three months to get = 36809.82*101.5% = $     37,361.97
Sell forward the dollars to get 37361.97*0.815 = €     30,450.00
Pay the euro loan of 30405 and take the profit of 30450-20405 = €             45.00
Profit in terms of US dollars = 45/0.815 = $             55.22